OpenStudy (anonymous):

x/y=3/4 , find 7x-4y/3x+y

OpenStudy (anonymous):

devide numerator and denominator of given equation by 'y' then put values of x/y in equation solve it

mathslover (mathslover):

\[\large{\frac{x}{y}=\frac{3}{4}}\] \[\large{x=\frac{3y}{4}}\] now put this in the eqn :7x-4y/3x+y

mathslover (mathslover):

\[\large{\frac{7x-4y}{3x+y}=\frac{\frac{21y}{4}-4y}{\frac{9y}{4}+y}}\]

mathslover (mathslover):

@9044614570 reply to me what you get?